The Everton College store paid $1650 for an order of 47 calculators. The store paid $8 for each scientific calculator. The others, all graphing calaulators, cost the store $57 each. How many of each type of calculator was ordered?
The store ordered ___ scientific calculators and ___ graphing calculators.

Thank you =)

Answer by rfer(16322)

x+y=47, x=47-y
8x+57y=1650
8(47-y)+57y=1650
376-8y+57y=1650
49y=1274
y=26
x=21
26*57+21*8=1650